a glass tube has mercury thread of length 100 cm at 0°C and 101.55cm at 100°C.if real expansion of mercury is 1.85*10^-4/C° .what is coefficient of linear expansion of glass.

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

The change in length of mercury in glass tube is given as

\Delta L=L_0\gamma_r\Delta T

\implies \gamma_r=0.55\times 10^{-4}/^{\circ}C
But\implies \gamma_r=\gamma_{Hg}-\gamma_{glass}
\implies \gamma_{glass}=1.3\times 10^{-4}/^{\circ}C=3\alpha_{glass}
\implies \alpha_{glass}=4.3\times 10^{-5}/^{\circ}C